North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog 7th to 13th January 2021

7th to 13th January 2021

A full week of activity to report -

7 January: Q6 - Ian Pearson and Bill Dobson met at lunchtime at Deviation Shed with the boiler in steam down at the MPD. However, Chris Kelly informed them just after lunch that he had had to fail the in house steam test due to the tea tray studs leaking. Once the boiler had cooled they would be removed. The insurance out of frames steam test arranged for the Monday was therefore cancelled and the boiler inspector informed.

12 January: Q6 - it was a cold frosty morning at Grosmont on Tuesday for Bill Dobson, Nigel Bill, Ian Pearson and Sean Bowler. Jon Bradley joined later in the afternoon. Gordon Wells also called in to rewire the kettle and water heater onto separate circuit breakers so they don’t trip out.

Bill removed the steam supply valve that feeds the vacuum brake. This was leaking steam and he found the seat required cutting. This job took most of the day, because, as is often the case, searching for cutters was a nightmare. In the end, one was borrowed from the MPD. A new brass washer is being made in place of a worn out one, so the valve will be re-fitted shortly. The water had been drained out of the boiler on Saturday morning. Nigel and Ian removed the pressure gauge and gauge glass protectors. In the afternoon, Sean completed the slide bar alignment then took both crossheads down to the MPD for measuring and checking. These will be re-metalled and machined. Later in the afternoon, with Jon assisting, the vacuum pipe was fitted.

An inspection of NELPG ladders was carried out, all of which are in Deviation Shed. One long, old, ladder has been scrapped due to defective rungs, and another has been labelled 'Not to be Used' as the integrated stabiliser has been bent and needs to be repaired. The footpath between Nos 8 and 7 road had become very cluttered with Items of the Q6 making the emergency door difficult to access. The items have been removed and the footpath is now clear.

Down at the MPD, tons of silt and ash was removed from drain sumps at the north end outside the MPD by a contract firm. We have been told that the water from Deviation Shed pit will have to be diverted away from the ash pit outside our building directly to the main drain to stop the ash getting into the traps and dumps.

13 January: Q6 - was another very cold but fine day with Bill Dobson, Ed Bolam, Ian Pearson and a late morning arrival of Nigel Bill who had picked up some consumables from Arco - 3 boxes of nitrile disposable gloves and 1 box of disposable face masks.

After getting the stove fired up, Bill, Ed and Ian fitted the steam heat pipes. They still require some securing, but are in position and also out of the way of the footpath. When Nigel arrived he spent most of the day fitting split pins to the slide bar bolts on both sides. Bill, Ed and Ian fitted fitted the front control sands rod and the left to right sands operating rod which connects to the front control rod. Ed also cleaned the the gauge glass protectors and replaced a spring on one of the protectors. Then he fitted the cylinder cocks, although some more fine adjustments are required.

There was also some tidying up in the workshop. An empty red cabinet, which has been in the way upstairs, was brought down and now sits on the bench at the south end of the workshop with a variety of tools stored in it.

Down at the MPD, a train was marshalled comprising of a well wagon with the boiler of 44806 secured to it and a brake van and Class 25 ready to move to Newbridge yard on Thursday. The boiler is going Riley engineering.

The next Q6 working parties will be on Tuesday 19 January with Ian Pearson and Bill Dobson, and Wednesday 20 January with Ian Pearson, Bill Dobson and Nigel Bill.